### v2

# Breaking Abbe’s diffraction limit with harmonic deactivation microscopy

Breaking Abbe’s diffraction limit with harmonic deactivation microscopy
Nonlinear optical microscopy provides elegant means for label-free imaging of biological samples and condensed-matter systems. The widespread areas of application could even be increased if resolution was improved, which the famous Abbe diffraction limit currently restrains. Super-resolution techniques can break the diffraction limit but most rely on fluorescent labeling. This makes them incompatible with (sub-)femtosecond temporal resolution and applications that demand the absence of labeling. Here, we introduce harmonic deactivation microscopy (HADES) for breaking the diffraction limit in non-fluorescent samples. By controlling the harmonic generation process on the quantum level with a second donut-shaped pulse, we confine the third harmonic generation to three times below the original focus size of a scanning microscope. We demonstrate that resolution improvement by deactivation is more efficient for higher harmonic orders, and only limited by the maximum applicable deactivation-pulse fluence. This provides a route towards sub-100 nm resolution in a regular nonlinear microscope.
Introduction
Far-field optical microscopy has impacted the fields of medicine, biology, and biophysics (?) like hardly any other technique. Microscopy is used as a daily diagnostics tool as well as equipment for cutting-edge scientific advancements. Microscopy benefits from its compactness, ease of use, and the possibility to image 2D as well as 3D objects (?).
Third harmonic generation (THG) is used as a label-free microscopy (?) method to study cells (?, ?), photoresists (?) and 2D materials (?). For THG microscopy, an excitation pulse is focused into a material and generates harmonics. The high penetration depth of the near-infrared (NIR) excitation light and the sensitivity to changes in chemical composition are central advantages of THG compared to other optical contrast mechanisms. One challenge preventing an even more widespread application is the limited resolution. The reported resolution of THG microscopes is approximately 1 µm or slightly less (?, ?), which is far from the 200-nm resolution limit commonly stated in a confocal fluorescence microscope (?). This apparent limitation stems from the fact that in typical confocal scanning microscopes, the resolution is proportional to the focus size of the excitation laser of the emission process. This laser is typically in the infrared for harmonic microscopy but in the visible or UV range for fluorescent microscopy. To improve this resolution limit of harmonic microscopy, the emission has to be confined below the diffraction limit. Here we show that the unique optical properties of high-harmonic genreation from condensed matter enable sub-diffraction harmonic emission and thus super-resolution microscopy.
High-harmonic generation (HHG) is a laser-driven frequency upconversion process (?, ?), that converts an ultra-short laser pulse into a shorter-wavelength pulse at integer multiples of the fundamental carrier frequency. The discovery of solid–state HHG in 2010 (?) also rejuvenated the interest in harmonic generation of lower orders below the bandgap (?) and drove the development of a stringent microscopic theoretical framework (?, ?) that complements the phenomenological framework of perturbative non-linear optics. On the quantum level, HHG is governed by laser-driven electron dynamics and contains signatures of the dynamic electronic structure of the generation medium that can be atoms (?), molecules (?, ?) and solids (?, ?). Inversely, slight variations to these electronic dynamics can have a large impact on the efficiency of HHG (?). In a variety of solids, reversible deactivation of high-harmonics followed the photoexcitation of charge carriers (?, ?, ?, ?, ?, ?). Different physical processes are considered to contribute to the deactivation (?, ?) such as state-blocking and ground-state depletion, a decrease in decoherence time (?, ?, ?), and control of electron-hole recollisions in shaped driving waveforms (?). The modulation depth, or contrast, achieved can be close to 100%, demonstrating near-unity-contrast emission control. This intensity modulation is the key enabler of super-resolution HHG microscopy by spatially confining HHG with an optical control pulse.
We took inspiration from the super-resolution techniques in fluorescence microscopy, which can provide resolutions below 10 nm (?). These techniques can be divided into two groups. The first group uses wide-field illumination of the sample and achieves super-resolution through stochastic on-and-off switching of fluorescence combined with localization algorithms (?, ?, ?, ?). The second group consists of deterministic methods that rely on targeted modulation or saturation of fluorescence (?, ?, ?, ?).
Here, we utilize the high modulation depth of HHG (?) for an approach similar to the one used in stimulated emission depletion (STED) microscopy (?) for HHG microscopy. We label this method as harmonic deactivation microscopy (HADES), which relies on an increase in incoherent scattering and disruption of the recombination process through a control beam. In the present manuscript, we focus on THG, but the method will work similarly and with even better resolution for higher harmonic orders. In HADES, we add a donut-shaped deactivation pulse to the infrared driving pulse of HHG. This donut pulse deactivates harmonic emission in the outer parts of the point-spread function (PSF) of the excitation pulse leading to a narrower PSF compared to THG. A similar methodology has been shown for linear microscopy in Silicon (?) and Graphene (?), relying on the plasma dispersion effect and absorption saturation, respectively. The latter herby was able to achieve a resolution of below 200 nm. For nonlinear microscopy utilizing coherent anti-Raman scattering methods were patented (?) and later shown experimentally (?). Improvements of the resolution of second-harmonic generation microscopy have been shown previously through coherent rescanning (?) and exploiting the nonlinear interaction in a structured illumination microscope (?). A two-fold resolution improvement is achieved through controlling the polarization state over the focus spot (?). In this work we demonstrate a two-dimensional resolution improvement utilizing increased incoherent scattering and recombination during the HHG process.
The combination of ultrafast temporal and nanoscale resolution will be a unique advantage of HADES: Due to the necessity of intramolecular vibrational energy redistribution before stimulated emission in STED, fluorescence super-resolution is inherently incompatible with femtosecond time resolution (?).
While deactivation is a general phenomenon, we chose NbO2 as a sample in the present manuscript. NbO2 is a strongly correlated material that undergoes a non-thermal ultrafast insulator-to-metal phase transition (IMT) upon photoexcitation. This phase transition makes strongly correlated materials interesting for a new class of electronic devices such as memristors (?, ?). In recent work HHG was used to probe this transition in NbO2 and measure its transition time scale to be 100 fs (?). The ability to add super-resolution to resolving IMTs will enable resolving the emergence of different phases with femtosecond temporal and nanometer spatial resolution.
In this letter, we show the time-dependent deactivation behavior of THG in NbO2 and how it’s governed by the fluence of the deactivation beam. These results are used to predict a change in the PSF of the third harmonic in HADES. We then change in our experiment the deactivation beam profile to a donut shape and confirm the narrower PSF. Conclusively, we improve the resolution of a scanning THG microscope by implementing HADES.
Results
Deactivation of third harmonic generation.
The concept of the experiment is illustrated in Fig. 1.
In regular harmonic microscopy, a NIR excitation laser generates harmonic multiples of its fundamental frequency.
For a Gaussian intensity distribution INIR of the fundamental NIR excitation laser in focus (Fig. 1a), the intensity distribution of the third harmonic will also be Gaussian (Fig. 1c). The spatial emission intensity distribution, called point-spread function (PSF), of HHG is narrower than the excitation pulse PSF and scales in the perturbative regime with I I, where is the power law exponent of the process which is approximately three for THG. In this work, we further deactivate HHG spatially by adding a pulse that is donut-shaped in focus to HHG (Fig. 1b). This donut-shaped pulse is achieved by imparting an orbital angular momentum onto the deactivation pulse by means of a spiral phase plate inserted in the far-field pulse profile. This concept allows deactivating HHG to a narrower emission peak below the diffraction limit (Fig. 1d) since HHG is deactivated in regions of non-zero intensity of the donut-shaped deactivation pulse. By narrowing the spot size below the diffraction limit, we can implement a super-resolution label-free scanning microscope.
We use a modified Mach-Zehnder interferometer to realize these experiments (Fig. S1). The fundamental HHG driving pulse is centered at a wavelength of 1830 nm, while the donut pulse is centered at 400 nm. Both pulses are linearly and parallel polarized. Detailed descriptions of all experiments can be found in the materials and methods section.
We first characterize the deactivation behavior by measuring the spectra of the emitted harmonics. We vary the delay between the pulses (Fig. 2a) as well as the intensity of the deactivation pulse (Fig. 2b). The spectral shape of the third harmonic (inset Fig. 2b) stays identical for all fluences. However, with increasing modulation fluence the amplitude and the integrated signal reduce. We normalize the integrated signal to its deactivation, i.e. a normalized deactivation of one means full deactivation of HHG so that there is no measurable signal, while zero means the signal is unchanged. When the driving pulse arrives before the deactivation pulse no change in the THG can be observed. This is followed by a sharp increase in the maximum deactivation when both pulses arrive simultaneously. At increasing time delays the signal recovers with a fast and slow exponential component, consistent with the results in (?).
The deactivation becomes stronger the higher the deactivation-pulse fluence. However, no dependence on the relative polarization angle could be observed during the experiments. At the highest deactivation-pulse fluence of 24 mJ/cm2, the integrated signal amounts to 2 of the reference at the temporal overlap of the two pulses. From this fluence dependence of the normalized deactivation, we determine a saturation fluence , defined as the fluence at half deactivation, of 2.5 mJ/cm2, while the intensity modulation depth is 98%. Comparing these results to deactivation curves for the fifth harmonic (grey curve Fig. 2b) shows that higher-order harmonics have a lower saturation fluence of 1.8 mJ/cm2.
While other STED-like techniques have been proposed and implemented in the past (?, ?), our approach - HADES - sets itself apart, both in its physical mechanism and its prospect for broad applicability. These intertwined features - broad applicability through a new mechanism - rest on how solid HHG can be deactivated by a modulation pulse, which is first and foremost a property of the HHG process on the quantum level as well as the driving field, but not of the material. The mechanistic understanding of solid harmonic generation has been rejuvenated in recent years (?) and extended beyond the phenomenological perturbative nonlinear picture. Solid HHG is understood two originate from two contributions following multiphoton excitation across the band gap. HHG is emitted from an intraband carrier current, as well as an interband polarization due to coherent electron-hole pairs. This interband polarization can be thought to originate from an electron-hole recollision. If a second modulation pulse is applied to control the HHG process, several phenomena occur as detailed in our recent perspective article on the topic (?) and briefly reviewed here: First, two-color waveforms strongly modulate HHG (?, ?), and the waveform of two overlapping electric fields (i.e at a time delay of zero) with incommensurate frequencies can decrease the overall intensity within a harmonic order (?). Second, right after excitation and for later times (i.e. for delays of and larger than zero) the coherent electron-hole pairs excited from the HHG driving pulse dephase faster due to the excited carrier population from the first pulse.In general, an excited carrier population promotes both accelerated carrier-carrier and carrier-phonon scattering, thus leading to a rapid loss of electron-hole coherence (?, ?). It is this electron-hole coherence which is neccessary for efficient HHG, so an increased dephasing causes an amplitude reduction and thus deactivation of HHG. In the specific case of NbO2, photo-excitation triggers an ultrafast phase transition on a time scale of 100 fs (?) into a metallic phase, where electron scattering is generally larger and thus the effect of HHG suppression through loss of electron-hole coherence is further enhanced. Scattering and the associated dephasing-time reduction act as amplitude suppression (?) window onto the time-domain emission of HHG, thus reducing the overall emission intensity. Both effects - waveform modulation and increased electron dephasing are contributing to deactivation in our experiment. This is evidenced by the observation of the highest levels of HHG deactivation at time zero for all pump fluences, which can thus be partially attributed to waveform modulation. Furthermore, the persisting high level of HHG deactivation after time zero can be attributed to increased scattering due to excited electrons and also an insulator-to-metal phase transition at higher pump fluences. At a time delay of zero, where the experiments described in this article are performed, both effects can thus play a role. Critically, none of these processes rely on resonant transitions in the material itself as the harmonic deactivation has been shown also for below bandgap pumping (?, ?, ?).
Point-spread function reduction.
We first estimate the expected resolution improvement of HADES for a one-dimensional model. We combine the fluence-dependent deactivation with the spatial distribution of the intensity of the donut pulse and multiply this with an initial Gaussian distribution of the emitted harmonics. Assuming a maximum fluence of 25 mJ/cm2, this one-dimensional model results in a PSF reduction by a factor of three (Fig. S2). This requires that the donut-shaped pulse has a constant azimuthal fluence distribution and that the maximum of the donut coincides with the FWHM of the THG, since a larger donut pulse lessens the effective applied fluence and a smaller donut pulse produces side wings which lead to smearing.
In the above model, we chose to use a cos2 function for modeling the donut-shaped deactivation profile as previously done in (?) for STED microscopy. We, therefore, modify the resulting formula to be adapted for our measurements as (?):
The first term containing the wavelength of the fundamental excitation laser , nonlinear order (three or five for detection of the third or fifth harmonic, respectively, in the perturbative regime) and numerical aperture NA, is the diffraction limit after Abbe modified for harmonic microscopy. The second term gives the resolution improvement based on the saturation level . If we assume a peak fluence of 24 mJ/cm2 and the measured saturation fluence for the third harmonic the resolution improves by a factor of 3.3, which is close to our one-dimensional model. For fifth harmonic generation, we calculate a resolution improvement of 3.8. This shows pathways to improve the resolution even further by decreasing the saturation fluence, e.g. via optimizing the deactivation scheme or by using higher-order harmonics that deactivate at lower fluences.
We now demonstrate this all-optical control and shrinkage of the PSF. A lens with a focal length of 25 cm generates the third harmonic in the sample (Fig. 3a). A home-built bright-field microscope images the PSF in transmission. We calculate the diffraction limit of the lens that focuses the driving pulses as 40.71.2 µm, which is the first term in eq. (1). The unchanged third harmonic PSF (Fig. 3b) has an FWHM of 401.5 µm, which is in agreement with the theoretical diffraction limit. When the donut pulse (inset figure 3c) deactivates the third harmonic emission the area that emits light is drastically reduced to an FWHM of 140.6 µm, while the peak intensity stays similar. This yields a reduction of factor 2.90.15, which matches the prediction of the one-dimensional model. Based on the experimental data, we calculated a two-dimensional model, which also shows a PSF reduction although this is limited by the diffraction from the uneven sample surface (Fig S3). The emission spot produced via HADES is therefore roughly three times smaller than Abbe’s diffraction limit in eq. (1). This paves the way towards a feasible approach for label-free super-resolution scanning microscopy through HADES, which we demonstrate in the next section in a proof-of-principle experiment.
Microscopic imaging.
For rapid sampling, the fundamental pulse of 800 nm was used for deactivation since a similar deactivation behavior to 400 nm was observed. We used a lens with a focal length of 5 cm as the focusing objective, which also captured the back-reflected THG for imaging (Fig. 4a). In the detection pathway, a series of bandpass filters (BP) select a specific harmonic order. A tube lens then focuses the filtered light onto an avalanche photodiode (APD). The APD measures signal with low average powers on a shot-to-shot basis, enabling rapid scanning of the sample. A boxcar integrator filters out noise outside the 2% duty cycle. Figures 4b) and c) show the recorded pictures for THG microscopy and HADES, respectively. These pictures depict the transition from an area with laser-induced damage to an undamaged sample area. The large dark area (area with no signal) has no thin film left, while the area emitting a lot of signal (red area) still has thin film. We analyze the origin of the sharp intensity features in the discussion. In comparison, the HADES picture visualizes smaller details compared to THG microscopy. The super-resolution becomes visible when convolving the HADES image with a Gaussian. This fully reproduces the THG microscopy image (Fig. S4), which demonstrates the resolution improvement by HADES. The two profiles (Fig. 4d) exemplify this resolution improvement further. Along the dotted line in the THG microscope, a single peak is imaged while by using HADES two peaks become distinguishable. These two peaks are separated by 9.7 µm. The dotted line depicts a single peak at the very edge of the thin film. With THG microscopy the FWHM of this emitter is 15.6 µm, while, with HADES the FWHM reduces to 7.1 µm. These examples can be quantified further by the use of a Fourier ring correlation (FRC) (?) analysis. First, a mask was applied to the Fourier transform of the images to suppress scanning artifacts. The FRC shows a higher correlation at larger spatial frequencies for HADES (Fig. S5). We choose the 1/7-threshold (?), as it is a common threshold for resolution estimation in 2D. This threshold yields a resolution limit of 8.6 µm and 5.5 µm for THG and HADES, respectively. We show, therefore, that the resolution improves in two dimensions at the same time.
Discussion
We now turn to discussing signal formation in HADES. The image shows an area of the NbO2 thin film that was thermally damaged before the HADES experiments. In bright-field microscopy, the dark area is visible as well, indicating that material was ablated (Fig. S6a) when damage occurred. The bright spot (dashed profile in Fig. 4) is located at a pointy tip of the thin film to pure substrate transition seen in dark-field microscopy (Fig. S6b). A combination of multiple effects can be the reason for the enhancement THG signal enhancement at this bright spot. First, the third-order susceptibility of the sample can be changed locally, since the NbO2 thin film generates harmonics more efficiently than the pure substrate. Second, near-field confinements of the electric field of the fundamental in subwavelength nanostructures can enhance the HHG efficiency (?). Indeed, the presence of such sub-wavelength nanostructures becomes visible under a scanning electron microscope (Fig. S7). During experimenting, we generally observed an enhancement of the signal when scanning over an edge in the sample, indicating near-field enhancement of the fundamental field, which boosts the efficiency of HHG locally.
Our current experimental procedure was challenged by the available low repetition rate (2 kHz) / high-pulse energy (3.5 mJ) laser system. The high pulse energies required a strong attenuation of both the modulation and excitation pulse by effective optical densities up to 5, which made detectable signals weak due to the 2-kHz repetition rate. These requirements made optimization more challenging especially for shorter focal lengths lenses. Furthermore, scanning microscopy as shown in Fig. 4 will benefit from either high-speed scanning stages or ideally a galvo-mirror system that scans the position of the focus through the objective, neither of which was available for the current experiment. Implementing both - higher repetition rate lasers and high-speed scanning will make taking microscopy images in seconds possible. This will also allow focusing the HHG driving pulses with high-NA objectives instead of lenses (Fig. 4). Since we have demonstrated the mechanism for breaking the diffraction limit in harmonic microscopy in this manuscript, we now understand that we can quantitatively predict the resolution improvement by means of Eq. 1 and the harmonic deactivation curve (Fig. 2b).
The ideas presented can be extended to tighter focusing conditions emitting higher harmonics. The in-depth analysis of the deactivation process measured here and by (?) delivers a three-times resolution improvement of THG microscopy and a 3.8-times resolution improvement for a fifth harmonic generation microscope. Considering a near-infrared excitation at 1800 nm and a high NA objective of 1.35, the expected resolutions for the specific set of wavelengths, polarizations, and material (NbO2) chosen here are 120 nm for THG and 80 nm for the fifth harmonic generation. These estimations are made based on eq. (1) and the experimental findings done in this paper and by (?), which yields a saturation fluence of 1.8 mJ/cm2 for the fifth harmonic. The maximum applicable fluence was herby always low enough to not damage the sample itself, meaning that laser-induced damage ultimately limits this technique. Deepening the understanding of the deactivation process for harmonic generation in materials will lead to an even better resolution by either reducing the saturation fluence, i.e. finding ways to deactivate HHG more efficiently, or by extending the highest applicable peak fluence. This principle is ultimately not limited to the deactivation of HHG by photoexcited carriers, which limits the maximum applicable fluence due to photo-induced damage. Other deactivation methods are also suited for this, i.e. control of the electron-hole recombination with a combination of below band-gap driving wavelengths might also be a suitable pathway (?). Ultimately, resolution in HADES is only limited by the real-space excursion of electrons and holes by the driving laser pulse. This excursion underlies the microscopic currents that radiate HHG, and corresponds to a real-space displacement on the single-digit nanometer scale, the exact value of which depends on electronic structure and laser parameters. As a hypothetical deactivation below this limit would truncate HHG even from the inner region without a deactivation pulse, this excursion length scale sets the ultimate resolution limit for HADES.
Furthermore, other optical nonlinearities such as coherent anti-Stokes Raman scattering (?), stimulated Raman scattering (?), or sum frequency generation (?) can be used to achieve super-resolution with similar concepts. However, the extreme nonlinearity of HHG creates a maximum sensitivity to control pulses. This likely makes HHG and related high-order frequency mixing processes the ideal candidates for resolution improvement by deactivation in particular, and optical emission control and optical switching in general. The wide range of solid-state materials, which have shown harmonic deactivation (?), make HADES an interesting candidate for metrology on dielectric and semiconducting materials, as well as, an important research tool for material science. Purley field effect-driven harmonic deactivation allows even application in biological samples.
Materials and Methods
Materials
The sample investigated in this study is the polycrystalline thin film NbO2, which was grown on a c-plane sapphire substrate through reactive bias target ion pulse deposition (RBTIBD), and the more information of the material synthesis are available in previous literature (?, ?). The Raman spectrum and X-ray diffraction 2 scan are both utilized to confirm the high quality and the bct lattice structure of our sample. Any other oxide components (NbO or Nb2O5) have not been found. The thickness of our sample is 115 nm, measured by X-ray reflectivity, and its surface roughness is less than 1 nm. Moreover, the sapphire substrates are double-side polished, ensuring the possibility of transmission measurements.
Methods
Laser system and experimental setup.
A Ti:Sapphire laser amplifier (Solstice, Spectra-Physics) produces a laser pulse at a central wavelength of 790 nm and with a pulse duration below 72 fs. The repetition rate is set to 2 kHz while the average output power is 7 W. A pulse splitter (BS) sends the laser light into each of the arms. In the excitation arm, the light pumps an optical parametric amplifier (OPA, Topas prime), which emits NIR light centered at 1800 nm. The idler of the parametric process was used to generate the third harmonic since only the idler can reach the desired wavelength. In the modulation arm, the delay of the 800 nm pulse can be adjusted by a mechanical delay stage. This is called the modulating arm. Afterward, a spiral phase plate (PP) converts the intensity profile into a donut shape. Optionally, a second harmonic generation BBO can be introduced after the spiral phase plate to change the central wavelength to 400 nm. Variable attenuators control the average power in each arm independently, while telescopes set the pulse size. A dichroic mirror (DM) combines the two pulses. After this, an objective lens (Obj) focuses this pulse bundle onto the sample (S). The modulation arm uses a combination of a half-wave plate and polarizer for a first attenuation, a second attenuation is achieved by a variable neutral density filter wheel. The excitation arm is attenuated by a series of neutral density filters. A schematic of the experiment can be found in the supplementary information (Fig. S1).
Deactivation measurements.
For the deactivation measurements, we removed the spiral phase plate from the modulation arm. A telescope focuses the modulation pulse so that it has an FWHM of 200 µm while the excitation NIR pulse has an FWHM of 40 µm. A tube lens focuses the back-reflected harmonics into a fiber spectrometer (Ocean Insight, QE Pro), that resolves the spectral composition of the harmonics.
PSF imaging.
A microscope objective (Mitutoyo Plan Apo NIR B) collimates the light from the sample with a magnification of 20x and an NA of 0.4. Subsequently, a tube lens with a focal length of 20 cm images them onto a camera. A pair of Bandpass filters blocked the donut pulse for PSF imaging. They have a central wavelength of 600 nm and a FWHM of 40 nm. The donut pulse was attenuated by an ND filter with an OD of 4.0 to avoid permanent damage. We evaluated the FWHM of the PSF by binarizing the images at a relative intensity of 0.5. Then an ellipse was fitted to the mask created that way. The FWHM represents the average radius of the fitted ellipse.
Microscopy pictures.
For the microscopy images both arms had to be attenuated by an effective optical density of 5. Otherwise, the high pulse energy of the laser could damage the thin-film surface. An 800 nm pulse deactivated the third harmonic response. To make scanning images we used a lens with a focal length of 5 cm and detected the reflected third harmonic emission. A silicon-based analog APD (Thorlabs) produces the signal out of the third harmonic generation. A preamplifier amplified the voltage output of the APD before the pulsed signal was integrated using a boxcar integrator. The Voltage output was then digitized by a data acquisition card. A piezo stage moved the sample through the focused pulse from spot to spot.
The images have a pixel size of 1 µm in both directions and are 51 x 51 pixels large. The images were scanned line by line. For the THG microscopy at each pixel 25 laser pulses were measured and each line was repeated 3 times, while for HADES each line was repeated 12 times with 25 laser pulses per pixel to ensure a similar count of photons for each pixel. The microscopy pictures displayed are the averaged values of each measurement, normalized to the maximum and minimum values. For the profile generation, the images were smoothed using a median filter with a pixel radius of 1.5. The profiles were generated with the profile plotter of ImageJ. The profiles were again normalized and interpolated by a spline algorithm for measurements of the FWHM of the features and measurements of the peak separation.
Acknowledgments
Acknowledgments:
We thank Christoph Morscher for help with dark-field images. We thank Antonios Pelekanidis for helping with the step-phase plates. We thank Matthijs Velsink for help with the APD. We thank the AMOLF support department for providing and helping with the measurement software and hardware.
Funding:
This work was conducted at the Advanced Research Center for Nanolithography, a public-private partnership between the University of Amsterdam (UvA), Vrije Universiteit Amsterdam (VU), Rijksuniversiteit Groningen (RUG), the Netherlands Organization for Scientific Research (NWO), and the semiconductor equipment manufacturer ASML and was (partly) financed by ‘Toeslag voor Topconsortia voor Kennis en Innovatie (TKI)’ from the Dutch Ministry of Economic Affairs and Climate Policy. This manuscript is part of a project that has received funding from the European Research Council (ERC) under the European Union’s Horizon Europe research and innovation programme (grant agreement no. 101041819, ERC Starting Grant ANACONDA). Z.N., L.G., S.W. and P.M.K. acknowledge support from the Open Technology Programme (OTP) by NWO, grant no. 18703. The project is also part of the VIDI research programme HIMALAYA with project number VI.Vidi.223.133 financed by NWO.
Author contribution:
K.M., L.G., S.W., and P.M.K. envisioned and conceptualized the experiments. K.M. and P.M.K. carried out a preliminary experiment, and K.M., M.L.S.G., and P.E. carried out the main experiments. K.M. and Z.N. analyzed the data. L.G. and Z.N. provided and characterized the sample. P.M.K. supervised the work. K.M. and P.M.K. wrote a first version of the manuscript that was finalized by all authors.
Competing interest:
The authors declare that they have no competing interests.
Data availability:
All data needed to evaluate the conclusions in the paper are present in the paper and/or the Supplementary Materials.

### v3

@@
 Breaking Abbe’s diffraction limit with harmonic deactivation microscopy
-Nonlinear optical microscopy provides elegant means for label-free imaging of biological samples and condensed-matter systems.
-The widespread areas of application could even be increased if resolution was improved, which the famous Abbe diffraction limit currently restrains.
-Super-resolution techniques can break the diffraction limit but most rely on fluorescent labeling.
+Nonlinear optical microscopy provides elegant means for label-free imaging of biological samples and condensed matter systems.
+The widespread areas of application could even be increased if resolution was improved, which is currently limited by the famous Abbe diffraction limit.
+Super-resolution techniques can break the diffraction limit but rely on fluorescent labeling.
 This makes them incompatible with (sub-)femtosecond temporal resolution and applications that demand the absence of labeling.
 Here, we introduce harmonic deactivation microscopy (HADES) for breaking the diffraction limit in non-fluorescent samples.
-By controlling the harmonic generation process on the quantum level with a second donut-shaped pulse, we confine the third harmonic generation to three times below the original focus size of a scanning microscope.
+By controlling the harmonic generation process on the quantum level with a second donut-shaped pulse, we confine the third harmonic generation to three times below the original focus size and use this pulse for scanning microscopy.
 We demonstrate that resolution improvement by deactivation is more efficient for higher harmonic orders, and only limited by the maximum applicable deactivation-pulse fluence.
 This provides a route towards sub-100 nm resolution in a regular nonlinear microscope.
+The new capability of label-free super-resolution can find immediate applications in condensed matter physics, semiconductor metrology, and biomedical imaging.
 Introduction
-Far-field optical microscopy has impacted the fields of medicine, biology, and biophysics (?) like hardly any other technique.
+Far-field optical microscopy has impacted the fields of medicine, biology, and biophysics (?) like kaum any other technique.
 Microscopy is used as a daily diagnostics tool as well as equipment for cutting-edge scientific advancements.
@@
 In a variety of solids, reversible deactivation of high-harmonics followed the photoexcitation of charge carriers (?, ?, ?, ?, ?, ?).
-Different physical processes are considered to contribute to the deactivation (?, ?) such as state-blocking and ground-state depletion, a decrease in decoherence time (?, ?, ?), and control of electron-hole recollisions in shaped driving waveforms (?).
+Different physical processes are considered to contribute to the deactivation (?) such as state-blocking and ground-state depletion, a decrease in decoherence time (?, ?, ?), and control of electron-hole recollisions in shaped driving waveforms (?).
 The modulation depth, or contrast, achieved can be close to 100%, demonstrating near-unity-contrast emission control.
@@
 Here, we utilize the high modulation depth of HHG (?) for an approach similar to the one used in stimulated emission depletion (STED) microscopy (?) for HHG microscopy.
-We label this method as harmonic deactivation microscopy (HADES), which relies on an increase in incoherent scattering and disruption of the recombination process through a control beam.
+We label this new method as harmonic deactivation microscopy (HADES).
 In the present manuscript, we focus on THG, but the method will work similarly and with even better resolution for higher harmonic orders.
@@
 This donut pulse deactivates harmonic emission in the outer parts of the point-spread function (PSF) of the excitation pulse leading to a narrower PSF compared to THG.
-A similar methodology has been shown for linear microscopy in Silicon (?) and Graphene (?), relying on the plasma dispersion effect and absorption saturation, respectively.
-The latter herby was able to achieve a resolution of below 200 nm.
-For nonlinear microscopy utilizing coherent anti-Raman scattering methods were patented (?) and later shown experimentally (?).
-Improvements of the resolution of second-harmonic generation microscopy have been shown previously through coherent rescanning (?) and exploiting the nonlinear interaction in a structured illumination microscope (?).
-A two-fold resolution improvement is achieved through controlling the polarization state over the focus spot (?).
-In this work we demonstrate a two-dimensional resolution improvement utilizing increased incoherent scattering and recombination during the HHG process.
+A similar methodology has been shown (?, ?) for linear microscopy and patented (?) for coherent anti-Raman scattering microscopy.
 The combination of ultrafast temporal and nanoscale resolution will be a unique advantage of HADES: Due to the necessity of intramolecular vibrational energy redistribution before stimulated emission in STED, fluorescence super-resolution is inherently incompatible with femtosecond time resolution (?).
@@
 In recent work HHG was used to probe this transition in NbO2 and measure its transition time scale to be 100 fs (?).
-The ability to add super-resolution to resolving IMTs will enable resolving the emergence of different phases with femtosecond temporal and nanometer spatial resolution.
-In this letter, we show the time-dependent deactivation behavior of THG in NbO2 and how it’s governed by the fluence of the deactivation beam.
-These results are used to predict a change in the PSF of the third harmonic in HADES.
-We then change in our experiment the deactivation beam profile to a donut shape and confirm the narrower PSF.
-Conclusively, we improve the resolution of a scanning THG microscope by implementing HADES.
+The ability to add super-resolution to resolving IMTs will enable resolving the emergence of new phases with femtosecond temporal and nanometer spatial resolution.
 Results
@@
 1c).
-The spatial emission intensity distribution, called point-spread function (PSF), of HHG is narrower than the excitation pulse PSF and scales in the perturbative regime with I I, where is the power law exponent of the process which is approximately three for THG.
+The spatial emission intensity distribution, called point-spread function (PSF), of HHG is narrower than for the excitation pulse and scales in the perturbative regime with I I, where is the power law exponent of the process which is approximately three for THG.
 In this work, we further deactivate HHG spatially by adding a pulse that is donut-shaped in focus to HHG (Fig.
@@
 This concept allows deactivating HHG to a narrower emission peak below the diffraction limit (Fig.
-1d) since HHG is deactivated in regions of non-zero intensity of the donut-shaped deactivation pulse.
+1d), since HHG is deactivated in regions of non-zero intensity of the donut-shaped deactivation pulse.
 By narrowing the spot size below the diffraction limit, we can implement a super-resolution label-free scanning microscope.
@@
 Detailed descriptions of all experiments can be found in the materials and methods section.
-We first characterize the deactivation behavior by measuring the spectra of the emitted harmonics.
+We first measure the spectra of the emitted harmonics.
 We vary the delay between the pulses (Fig.
@@
 We normalize the integrated signal to its deactivation, i.e.
-a normalized deactivation of one means full deactivation of HHG so that there is no measurable signal, while zero means the signal is unchanged.
+a normalized deactivation of one means that we measure no signal, while zero means the signal is unchanged.
 When the driving pulse arrives before the deactivation pulse no change in the THG can be observed.
@@
 The deactivation becomes stronger the higher the deactivation-pulse fluence.
-However, no dependence on the relative polarization angle could be observed during the experiments.
 At the highest deactivation-pulse fluence of 24 mJ/cm2, the integrated signal amounts to 2 of the reference at the temporal overlap of the two pulses.
@@
 2b) shows that higher-order harmonics have a lower saturation fluence of 1.8 mJ/cm2.
-While other STED-like techniques have been proposed and implemented in the past (?, ?), our approach - HADES - sets itself apart, both in its physical mechanism and its prospect for broad applicability.
-These intertwined features - broad applicability through a new mechanism - rest on how solid HHG can be deactivated by a modulation pulse, which is first and foremost a property of the HHG process on the quantum level as well as the driving field, but not of the material.
-The mechanistic understanding of solid harmonic generation has been rejuvenated in recent years (?) and extended beyond the phenomenological perturbative nonlinear picture.
-Solid HHG is understood two originate from two contributions following multiphoton excitation across the band gap.
-HHG is emitted from an intraband carrier current, as well as an interband polarization due to coherent electron-hole pairs.
-This interband polarization can be thought to originate from an electron-hole recollision.
-If a second modulation pulse is applied to control the HHG process, several phenomena occur as detailed in our recent perspective article on the topic (?) and briefly reviewed here: First, two-color waveforms strongly modulate HHG (?, ?), and the waveform of two overlapping electric fields (i.e at a time delay of zero) with incommensurate frequencies can decrease the overall intensity within a harmonic order (?).
+Suppression of HHG can be attributed to two processes.
+First, two-color waveforms strongly modulate HHG (?, ?), and the waveform of two overlapping electric fields (i.e at a time delay of zero) with incommensurate frequencies can decrease the overall intensity within a harmonic order (?).
 Second, right after excitation and for later times (i.e.
-for delays of and larger than zero) the coherent electron-hole pairs excited from the HHG driving pulse dephase faster due to the excited carrier population from the first pulse.In general, an excited carrier population promotes both accelerated carrier-carrier and carrier-phonon scattering, thus leading to a rapid loss of electron-hole coherence (?, ?).
-It is this electron-hole coherence which is neccessary for efficient HHG, so an increased dephasing causes an amplitude reduction and thus deactivation of HHG.
-In the specific case of NbO2, photo-excitation triggers an ultrafast phase transition on a time scale of 100 fs (?) into a metallic phase, where electron scattering is generally larger and thus the effect of HHG suppression through loss of electron-hole coherence is further enhanced.
+for delays of and larger than zero) electron scattering is increased.
+On the one hand, an excited carrier fraction increases scattering and reduces the dephasing time (?, ?), i.e.
+the time until which electron-hole pairs recombine coherently (?).
+On the other hand, excitation of NbO2 triggers an ultrafast phase transition on a time scale of 100 fs (?) into a metallic phase, where electron scattering is generally larger.
 Scattering and the associated dephasing-time reduction act as amplitude suppression (?) window onto the time-domain emission of HHG, thus reducing the overall emission intensity.
 Both effects - waveform modulation and increased electron dephasing are contributing to deactivation in our experiment.
-This is evidenced by the observation of the highest levels of HHG deactivation at time zero for all pump fluences, which can thus be partially attributed to waveform modulation.
+This is evidenced by the observation of highest levels of HHG deactivation at time zero for all pump fluences, which can thus be partially attributed to waveform modulation.
 Furthermore, the persisting high level of HHG deactivation after time zero can be attributed to increased scattering due to excited electrons and also an insulator-to-metal phase transition at higher pump fluences.
 At a time delay of zero, where the experiments described in this article are performed, both effects can thus play a role.
-Critically, none of these processes rely on resonant transitions in the material itself as the harmonic deactivation has been shown also for below bandgap pumping (?, ?, ?).
 Point-spread function reduction.
-We first estimate the expected resolution improvement of HADES for a one-dimensional model.
+We first estimate the expected resolution improvement of HADES numerically.
 We combine the fluence-dependent deactivation with the spatial distribution of the intensity of the donut pulse and multiply this with an initial Gaussian distribution of the emitted harmonics.
-Assuming a maximum fluence of 25 mJ/cm2, this one-dimensional model results in a PSF reduction by a factor of three (Fig.
+Assuming a maximum fluence of 25 mJ/cm2, this numerical model results in a PSF reduction by a factor of three (Fig.
 S2).
-This requires that the donut-shaped pulse has a constant azimuthal fluence distribution and that the maximum of the donut coincides with the FWHM of the THG, since a larger donut pulse lessens the effective applied fluence and a smaller donut pulse produces side wings which lead to smearing.
-In the above model, we chose to use a cos2 function for modeling the donut-shaped deactivation profile as previously done in (?) for STED microscopy.
-We, therefore, modify the resulting formula to be adapted for our measurements as (?):
+This requires that the donut has a constant azimuthal fluence distribution and that the maximum of the donut coincides with the FWHM of the THG.
+We compare the result to a modified formula for the resolution limit of STED microscopy (?):
 The first term containing the wavelength of the fundamental excitation laser , nonlinear order (three or five for detection of the third or fifth harmonic, respectively, in the perturbative regime) and numerical aperture NA, is the diffraction limit after Abbe modified for harmonic microscopy.
 The second term gives the resolution improvement based on the saturation level .
-If we assume a peak fluence of 24 mJ/cm2 and the measured saturation fluence for the third harmonic the resolution improves by a factor of 3.3, which is close to our one-dimensional model.
+If we assume a peak fluence of 24 mJ/cm2 and the measured saturation fluence for the third harmonic the resolution improves by a factor of 3.3, which is close to our numerical simulations.
 For fifth harmonic generation, we calculate a resolution improvement of 3.8.
@@
 When the donut pulse (inset figure 3c) deactivates the third harmonic emission the area that emits light is drastically reduced to an FWHM of 140.6 µm, while the peak intensity stays similar.
-This yields a reduction of factor 2.90.15, which matches the prediction of the one-dimensional model.
-Based on the experimental data, we calculated a two-dimensional model, which also shows a PSF reduction although this is limited by the diffraction from the uneven sample surface (Fig S3).
-The emission spot produced via HADES is therefore roughly three times smaller than Abbe’s diffraction limit in eq.
+This yields a reduction of factor 2.90.15, which matches the prediction of the numerical model.
+HADES’s spot therefore is roughly three times smaller than Abbe’s diffraction limit in eq.
 (1).
@@
 Microscopic imaging.
-For rapid sampling, the fundamental pulse of 800 nm was used for deactivation since a similar deactivation behavior to 400 nm was observed.
+For rapid sampling, the fundamental pulse of 800 nm was used since a similar deactivation behavior to 400 nm was observed.
 We used a lens with a focal length of 5 cm as the focusing objective, which also captured the back-reflected THG for imaging (Fig.
@@
 These pictures depict the transition from an area with laser-induced damage to an undamaged sample area.
-The large dark area (area with no signal) has no thin film left, while the area emitting a lot of signal (red area) still has thin film.
-We analyze the origin of the sharp intensity features in the discussion.
 In comparison, the HADES picture visualizes smaller details compared to THG microscopy.
@@
 This fully reproduces the THG microscopy image (Fig.
-S4), which demonstrates the resolution improvement by HADES.
+S3), whch demonstrates the resolution improvemnet by HADES.
 The two profiles (Fig.
-4d) exemplify this resolution improvement further.
+4d) quantify this resolution improvement further.
 Along the dotted line in the THG microscope, a single peak is imaged while by using HADES two peaks become distinguishable.
@@
 With THG microscopy the FWHM of this emitter is 15.6 µm, while, with HADES the FWHM reduces to 7.1 µm.
-These examples can be quantified further by the use of a Fourier ring correlation (FRC) (?) analysis.
-First, a mask was applied to the Fourier transform of the images to suppress scanning artifacts.
-The FRC shows a higher correlation at larger spatial frequencies for HADES (Fig.
-S5).
-We choose the 1/7-threshold (?), as it is a common threshold for resolution estimation in 2D.
-This threshold yields a resolution limit of 8.6 µm and 5.5 µm for THG and HADES, respectively.
+Notably, the two profiles are close to perpendicular to each other.
 We show, therefore, that the resolution improves in two dimensions at the same time.
@@
 We now turn to discussing signal formation in HADES.
-The image shows an area of the NbO2 thin film that was thermally damaged before the HADES experiments.
+The image in Fig.
+4 shows an area of the NbO2 thin film that was thermally damaged prior to the HADES experiments.
 In bright-field microscopy, the dark area is visible as well, indicating that material was ablated (Fig.
-S6a) when damage occurred.
+S4a) when damage occured.
 The bright spot (dashed profile in Fig.
 4) is located at a pointy tip of the thin film to pure substrate transition seen in dark-field microscopy (Fig.
-S6b).
-A combination of multiple effects can be the reason for the enhancement THG signal enhancement at this bright spot.
+S4b).
+A combination of multiple effects can be the reason for the enhancement THG signal enhancement at htis bright spot.
 First, the third-order susceptibility of the sample can be changed locally, since the NbO2 thin film generates harmonics more efficiently than the pure substrate.
@@
 Indeed, the presence of such sub-wavelength nanostructures becomes visible under a scanning electron microscope (Fig.
-S7).
+S5).
 During experimenting, we generally observed an enhancement of the signal when scanning over an edge in the sample, indicating near-field enhancement of the fundamental field, which boosts the efficiency of HHG locally.
@@
 2b).
-The ideas presented can be extended to tighter focusing conditions emitting higher harmonics.
-The in-depth analysis of the deactivation process measured here and by (?) delivers a three-times resolution improvement of THG microscopy and a 3.8-times resolution improvement for a fifth harmonic generation microscope.
-Considering a near-infrared excitation at 1800 nm and a high NA objective of 1.35, the expected resolutions for the specific set of wavelengths, polarizations, and material (NbO2) chosen here are 120 nm for THG and 80 nm for the fifth harmonic generation.
-These estimations are made based on eq.
-(1) and the experimental findings done in this paper and by (?), which yields a saturation fluence of 1.8 mJ/cm2 for the fifth harmonic.
-The maximum applicable fluence was herby always low enough to not damage the sample itself, meaning that laser-induced damage ultimately limits this technique.
+The in-depth analysis of the deactivation process measured here and by (?) delivers a three-time resolution improvement of THG microscopy and a 3.8-time resolution improvement for a fifth harmonic generation microscope.
+In conjunction with a near-infrared excitation at 1800 nm and a high NA objective of 1.35, the expected resolutions for the specific set of wavelengths, polarizations, and material (NbO2) chosen here are 120 nm for THG and 80 nm for the fifth harmonic generation.
 Deepening the understanding of the deactivation process for harmonic generation in materials will lead to an even better resolution by either reducing the saturation fluence, i.e.
@@
 This excursion underlies the microscopic currents that radiate HHG, and corresponds to a real-space displacement on the single-digit nanometer scale, the exact value of which depends on electronic structure and laser parameters.
-As a hypothetical deactivation below this limit would truncate HHG even from the inner region without a deactivation pulse, this excursion length scale sets the ultimate resolution limit for HADES.
+As a hypothetical deactivation below this limit would actually truncate HHG even from the inner region without a deactivation pulse, this excursion length scale sets the ultimate resolution limit for HADES.
 Furthermore, other optical nonlinearities such as coherent anti-Stokes Raman scattering (?), stimulated Raman scattering (?), or sum frequency generation (?) can be used to achieve super-resolution with similar concepts.
@@
 This likely makes HHG and related high-order frequency mixing processes the ideal candidates for resolution improvement by deactivation in particular, and optical emission control and optical switching in general.
-The wide range of solid-state materials, which have shown harmonic deactivation (?), make HADES an interesting candidate for metrology on dielectric and semiconducting materials, as well as, an important research tool for material science.
-Purley field effect-driven harmonic deactivation allows even application in biological samples.
 Materials and Methods
@@
 This is called the modulating arm.
-Afterward, a spiral phase plate (PP) converts the intensity profile into a donut shape.
-Optionally, a second harmonic generation BBO can be introduced after the spiral phase plate to change the central wavelength to 400 nm.
+Afterwards, a vortex phase plate (PP) converts the intensity profile into a donut shape.
+Optionally, a second harmonic generation BBO can be introduced after the vortex phase plate to change the central wavelength to 400 nm.
 Variable attenuators control the average power in each arm independently, while telescopes set the pulse size.
@@
 Deactivation measurements.
-For the deactivation measurements, we removed the spiral phase plate from the modulation arm.
+For the deactivation measurements, we removed the vortex phase plate from the modulation arm.
 A telescope focuses the modulation pulse so that it has an FWHM of 200 µm while the excitation NIR pulse has an FWHM of 40 µm.
@@
 The donut pulse was attenuated by an ND filter with an OD of 4.0 to avoid permanent damage.
+The images were taken
 We evaluated the FWHM of the PSF by binarizing the images at a relative intensity of 0.5.
@@
 The images were scanned line by line.
-For the THG microscopy at each pixel 25 laser pulses were measured and each line was repeated 3 times, while for HADES each line was repeated 12 times with 25 laser pulses per pixel to ensure a similar count of photons for each pixel.
+For the THG microscopy at each pixel 25 laser pulses were measured and each line was repeated 3 times, while for HADES each line was repeated 12 times with 25 laser pulses per pixel.
 The microscopy pictures displayed are the averaged values of each measurement, normalized to the maximum and minimum values.
@@
 Acknowledgments
-Acknowledgments:
-We thank Christoph Morscher for help with dark-field images.
-We thank Antonios Pelekanidis for helping with the step-phase plates.
-We thank Matthijs Velsink for help with the APD.
-We thank the AMOLF support department for providing and helping with the measurement software and hardware.
 Funding: